### v1

# Kendall’s $\tau$ Coefficient for Logits Distillation

Abstract
Knowledge distillation typically employs the Kullback-Leibler (KL) divergence to constrain the student model’s output to match the soft labels provided by the teacher model exactly. However, sometimes the optimization direction of the KL divergence loss is not always aligned with the task loss, where a smaller KL divergence could lead to erroneous predictions that diverge from the soft labels. This limitation often results in suboptimal optimization for the student. Moreover, even under temperature scaling, the KL divergence loss function tends to overly focus on the larger-valued channels in the logits, disregarding the rich inter-class information provided by the multitude of smaller-valued channels. This hard constraint proves too challenging for lightweight students, hindering further knowledge distillation. To address this issue, we propose a plug-and-play ranking loss based on Kendall’s $\tau$ coefficient, called Rank-Kendall Knowledge Distillation (RKKD). RKKD balances the attention to smaller-valued channels by constraining the order of channel values in student logits, providing more inter-class relational information. The rank constraint on the top-valued channels helps avoid suboptimal traps during optimization. We also discuss different differentiable forms of Kendall’s $\tau$ coefficient and demonstrate that the proposed ranking loss function shares a consistent optimization objective with the KL divergence. Extensive experiments on the CIFAR-100 and ImageNet datasets show that our RKKD can enhance the performance of various knowledge distillation baselines and offer broad improvements across multiple teacher-student architecture combinations.
1 Introduction
The recent advancements in deep neural networks (DNN) have significantly enhanced performance in the field of computer vision. This progress is driven by the pursuit of higher network capabilities, leading to an ongoing deepening of visual models and a continuous increase in the number of parameters. However, the heightened computational and storage costs associated with complex networks have limited their applicability. To address this, knowledge distillation (KD) has been proposed to obtain performant lightweight models. Typically, knowledge distillation involves a well-trained heavy teacher model and an untrained lightweight student model. The same data is fed to both the teacher and the student, with the teacher’s outputs serving as soft labels for training the student. By constraining the student to produce predictions that match the soft labels, the knowledge in the teacher model is transferred to the lightweight student model.
Most logit-based KD methods adhere to the paradigm introduced by Hinton [12], employing the Kullback-Leibler (KL) divergence [17] to align the logits output by the student and teacher models. The optimization goal of KL divergence is to achieve identical outputs between the student and teacher, thereby transferring as much knowledge as possible from the teacher to the student. However, given the large and complex nature of the teacher model, matching the output of the teacher without additional assistance is too challenging a task for the lightweight student. Recent studies [2; 27] have attempted to reduce the difficulty of the student’s learning task, such as reducing the gap between students and teachers by logits normalization, yet methods to fully leverage the information contained within logits remain underexplored.
Figure 1: Logits Value Distributions. (a) The original logits output by teacher and student. (b) The logits of teacher and student after a softmax with $T=4$.
Figure 2: Suboptimal Case of the KL Divergence.
We notice that the optimization direction of KL divergence sometimes conflicts with the task loss optimization direction. As illustrated in Fig. 2, compared to Student 2, Student 1 exhibits a smaller KL divergence with the teacher; however, Student 2 achieves the correct classification result consistent with the teacher, while Student 1 does not. Therefore, a smaller KL divergence does not always lead to correct classification outcomes. On the other hand, the KL divergence places significant emphasis on the large channels, often at the expense of the smaller ones. In practice, however, the smaller-valued channels in logits constitute a substantial proportion of the teacher and student outputs on the training set (as shown in Fig. 1(a)). Although a temperature parameter can be added to the softmax equation to control the sharpness of the logits distribution, we still find this imbalance exists after softmax (as shown in Fig. 1(b)). We posit that these smaller-valued channels are equally rich in teacher knowledge, such as inter-class relationships, and that effectively utilizing this information is beneficial for the student’s learning.
In this paper, we introduce a plug-and-play Ranking Loss based on Kendall’s $\tau$ Coefficient [15] to assist the optimization process of KL divergence, called Rank-Kendall Knowledge Distillation (RKKD). This rank loss equitably focuses on the consistency of the magnitude order of values within logits, supplementing the student’s optimization process with knowledge from the smaller-valued channels and information about inter-class relationships in predictions. Moreover, the correct classification predicted by the teacher invariably possesses the top $1$st logit value, and the ranking loss constrains the correct category to the foremost rank, aiding the KL divergence in correcting the optimization direction to align with task loss. Specifically, we first explore three differentiable forms of Kendall’s $\tau$ coefficient and design the ranking loss function based on these. Subsequently, we demonstrate that the optimal optimization objective of the proposed ranking loss encompasses that of KL divergence, indicating that RKKD can assist the optimization process. By adding the ranking loss as an auxiliary loss to the existing knowledge distillation baseline, it can enhance performance for various teacher-student pairings in distillation tasks, suggesting that RKKD can serve as a plug-and-play method to improve the performance of knowledge distillation approaches. In conclusion, our contributions are as follows:
  • To mitigate the issues arising from the overly direct nature of KL divergence, which leads to the disregard of smaller-valued channels and suboptimal optimization directions, we introduce a plug-and-play ranking loss function to assist the optimization process in knowledge distillation (Rank-Kendall Knowledge Distillation, RKKD). Drawing on Kendall’s $\tau$ coefficient, we explore and discuss three different versions of the ranking loss function to improve the efficacy of knowledge distillation.
  • We establish that the proposed ranking loss function aligns with the optimal optimization objectives of KL divergence, and we substantiate this alignment through visualizations and empirical evidence, demonstrating the benefits of the ranking loss function for the optimization of distillation tasks.
  • Extensive experiments were conducted on a variety of teacher-student architectures using the CIFAR-100 and ImageNet datasets. Our findings confirm the widespread effectiveness of the proposed ranking loss in distillation tasks, and its role as a plug-and-play auxiliary function provides substantial support for the training of distillation tasks.
2 Related Works
Knowledge Distillation. Knowledge distillation, initially proposed by Hinton [12], serves as a method for model compression and acceleration, aiming to transfer knowledge from a heavy teacher model to a lightweight student model. By feeding the same samples, the teacher can produce soft labels, and training the student with these soft labels allows the transfer of knowledge from the teacher model to the student. Knowledge distillation tasks can be mainly divided into two categories: feature-based distillation [3; 40; 36; 8; 21] and logit-based distillation [12; 20; 42; 14; 27; 5; 34]. Feature-based distillation additionally utilizes the model’s intermediate features, providing more information to the student and enabling the student to learn at the feature level from the teacher, often resulting in better performance. Considering safety and privacy, the intermediate outputs of the model are often not obtainable; hence, logit-based methods that only use the model outputs for distillation offer better versatility and robustness. Our proposed method can act as a plug-and-play module added to logit-based methods, offering higher flexibility and further enhancing the performance of logit-based methods.
Recent knowledge distillation methods have found that overly strict constraints can sometimes hinder the student’s transfer of knowledge from the teacher’s soft labels. For instance, [2] discovered that differences in feature sizes in feature-based methods could limit the student’s learning; while [27] found that using the same temperature for both teacher and student in logit-based methods could affect further improvements in student performance. To reduce the learning difficulty for the student, [27; 2] provided methods to ease the logit matching difficulty between student and teacher, yet we find that methods providing additional guidance to the student have not been fully explored.
Ranking Loss in Knowledge Distillation. For knowledge distillation, ranking loss is a relaxed constraint that can provide rich inter-class information. It was first applied in the distillation of recommendation systems [23; 29; 6; 22; 35], [18] explored the application of ranking loss in object detection tasks, and [7] discussed the role of ranking in the distillation of language model tasks. However, the exploration of ranking loss in logit-based image classification task distillation is not yet comprehensive. We find that the KL divergence used for distillation in classification tasks is an overly strict constraint that tends to overlook information from smaller-valued channels and may lead to suboptimal results. Our proposed method leverages ranking loss to balance the model’s attention to larger and smaller-valued channels, while also using inter-class relationships to help the model avoid suboptimal outcomes.
3 Preliminary
Most logit-based distillation methods adhere to the original KD proposed by Hinton [12], which transfers knowledge by matching the logit outputs of the student and teacher. This setting is more generalizable, allowing for distillation solely through outputs when the internal structures of the student and teacher are invisible. For a given dataset $\mathcal{D}$, assuming there are $C$ categories and $N$ samples, we possess a teacher model $f_{t}$ and a student model $f_{s}$. For a given sample $\mathcal{I}\in\mathcal{D}$, we can obtain the outputs of the teacher and student model, denoted as $z^{t}=f_{t}(\mathcal{I}),z^{s}=f_{s}(\mathcal{I})$ where $z^{t},z^{s}\in\mathbb{R}^{1\times C}$. Through a softmax function with temperature, the outputs are processed into the prediction vectors $q_{t},q_{s}\in\mathbb{R}^{1\times C}$ finally:
$q^{t}_{i}=\frac{exp(z^{t}_{i}/\mathcal{T})}{\sum_{j=1}^{C}exp(z^{t}_{j}/\mathcal{T})}$ (1)
$q^{t}_{i}=\frac{exp(z^{s}_{i}/\mathcal{T})}{\sum_{j=1}^{C}exp(z^{s}_{j}/\mathcal{T})}$ (2)
where $\mathcal{T}$ is a temperature parameter, $z_{i}$ represents the logit value of the $i$-th channel of the model output, $q_{i}$ represents the predicted probability for the target being the $i$-th class. The KL divergence loss function used to constrain the student and teacher logits is of the following form:
$\mathcal{L}_{KL}(q^{t},q^{s})=\sum_{i=1}^{C}q_{i}^{t}\cdot log(\frac{q_{i}^{t}}{q_{i}^{s}})$ (3)
It can be observed that in Eq. 3, the importance of the match between the student and teacher at the $i$-th channel is influenced by the coefficient $q_{i}^{t}$. This implies that channels with smaller logit values receive less attention, leading to the KL divergence’s disregard for smaller channels.
4 Ranking Loss Based on Kendall’s $\tau$ Coefficient
In this section, we introduce a ranking loss function designed to constrain the order of magnitude of channel values in the logits output by the student model. With the aid of the ranking loss, the knowledge distillation task can balance the overemphasis on larger logit values and the neglect of smaller ones as measured by the KL divergence. Additionally, the ranking loss imposes a constraint on the leading channel value, which helps to correct the optimization direction and avoid suboptimal solutions. In Section4.1, we will introduce the ranking loss function and employ Kendall’s $\tau$ coefficient to compute the ordinal consistency between the teacher and student logits. Section 4.2 will discuss the differentiable form of Kendall’s $\tau$ coefficient and, based on this, design three distinct forms of ranking loss functions. In Section 4.3, we will demonstrate that the optimization objectives of Rank Loss and KL divergence are aligned, and that Rank Loss can eliminate some of the suboptimal aspects of KL divergence, benefiting the distillation process for the student.
4.1 Differentiable Kendall’s $\tau$ Coefficient
In order to make the loss function pay more attention to the information provided by smaller channels as well as to help correct the suboptimal problem in Figure 1, we introduce the ranking loss based on Kendall’s $\tau$ coefficient. By pairing each of the $C$ channels, we can obtain $\frac{C(C-1)}{2}$ pairs, whose Kendall’s $\tau$ coefficient is expressed as follows:
$\tau=\frac{P_{c}-P_{d}}{\frac{1}{2}C(C-1)}$ (4)
where $P_{c}$ represents concordant pairs and $P_{d}$ represents discordant pairs. Kendall’s $\tau$ coefficient provides an expression of ordinal similarity. For the teacher and student logits, a channel pair $(i,j)$ is considered concordant if the signs of $(z^{t}_{i}-z^{t}_{j})$ for the teacher and$(z^{s}_{i}-z^{s}_{j})$ for the student are the same; otherwise, the pair is discordant. We substitute the logits of the teacher and student in the following manner:
$\tau=\frac{\sum_{i}\sum_{j}sgn(z^{t}_{i}-z^{t}_{j})\cdot sgn(z^{s}_{i}-z^{s}_{j})}{\frac{1}{2}C(C-1)}$ (5)
where $sgn()$ represents sign function. While we have quantified the ordinal similarity between the teacher and student logits, the aforementioned formula is non-smooth. To utilize the ordinal similarity for gradient computation, we approximate the sign function with the $tanh$ function to convert it into a differentiable form:
$\tau_{d}$ $=\frac{\sum_{i}\sum_{j}tanh(z^{t}_{i}-z^{t}_{j})\cdot tanh(z^{s}_{i}-z^{s}_{j})}{\frac{1}{2}C(C-1)}$ (6)
$=\frac{2}{C(C-1)}\cdot\sum_{i}\sum_{j}\left(1-\frac{2}{1+e^{(z^{t}_{i}-z^{t}_{j})\cdot k}}\right)\cdot\left(1-\frac{2}{1+e^{(z^{s}_{i}-z^{s}_{j})\cdot k}}\right)$
where $k$ is a parameter that controls the steepness of the function; a larger $k$ causes the $tanh$ function to more closely approximate the sign function. By negating the similarity measure in Eq. 6, it can serve as a loss function to enforce the consistency of the logits order between the teacher and student:
$\mathcal{L}_{RK}=-\tau_{d}=-\frac{2}{C(C-1)}\cdot\sum_{i}\sum_{j}\left(1-\frac{2}{1+e^{(z^{t}_{i}-z^{t}_{j})\cdot k}}\right)\cdot\left(1-\frac{2}{1+e^{(z^{s}_{i}-z^{s}_{j})\cdot k}}\right)$ (7)
The overall loss function is formulated as follows:
$\mathcal{L}=\alpha\mathcal{L}_{KL}+\beta\mathcal{L}_{CE}+\gamma\mathcal{L}_{RK}$ (8)
where $\mathcal{L}_{KL}$ denotes KL divergence loss, $\mathcal{L}_{CE}$ denotes Cross-Entropy loss, $\mathcal{L}_{RK}$ denotes our ranking loss, and $\alpha,\beta,\gamma$ are hyper-parameters.
4.2 Different Forms of Ranking Loss
In the initial application of the ranking loss[43], it is necessary to compute the gradients for two input vectors separately. However, In the scenario of distillation, the soft labels from the teacher model do not require gradients. Therefore, we propose three variants of the diff-Kendall ranking loss suitable for distillation scenarios, aimed at further exploring the role of ranking loss in distillation. From Eq.5, we can derive an equivalent form:
$\tau=\frac{\sum_{i}\sum_{j}sgn((z^{t}_{i}-z^{t}_{j})\cdot(z^{s}_{i}-z^{s}_{j}))}{\frac{1}{2}C(C-1)}$ (9)
For Eq.5 and Eq.9, we can transform them into a differential form of ranking loss, and gradients are not computed for the output part corresponding to the teacher, which is:
$L_{\tau}^{form1}=\frac{\sum_{i}\sum_{j}tanh(z^{t}_{i}-z^{t}_{j})_{detach}\cdot tanh(z^{s}_{i}-z^{s}_{j})}{\frac{1}{2}C(C-1)}$ (10)
$L_{\tau}^{form2}=\frac{\sum_{i}\sum_{j}tanh[(z^{t}_{i}-z^{t}_{j})_{detach}\cdot(z^{s}_{i}-z^{s}_{j})]}{\frac{1}{2}C(C-1)}$ (11)
Further, since the gradient of the teacher’s output is not required, we can directly use the sign function instead of the tanh function to obtain the ranking loss:
$L_{\tau}^{form3}=\frac{\sum_{i}\sum_{j}sgn(z^{t}_{i}-z^{t}_{j})_{detach}\cdot tanh(z^{s}_{i}-z^{s}_{j})}{\frac{1}{2}C(C-1)}$ (12)
4.3 How RKKD Benefit Optimizing Logits Distillation
In this section, we analyze the optimal solution domain for Kullback-Leibler (KL) divergence and Ranking loss to show how ranking loss affects the optimization process of KL divergence. For KL divergence, the optimal solution domain is achieved when the distribution of the student model aligns perfectly with that of the teacher model, which is equivalent to a linear mapping on the logits.:
$p_{i}^{t}=p_{i}^{s}\quad\forall i\quad\Rightarrow\quad z^{t}_{i}=z^{s}_{i}+c\quad\forall i$ (13)
$\ p_{i}^{t}$ and $\ p_{i}^{s}$ respectively represent the probability outputs of the teacher and the student for the i-th class, and $\ c$ represents any constant number. For the ranking loss, the optimal solution domain is such that for any two indices, the order of the logits output by the teacher and the student is consistent:
$\hskip 5.3ptsgn(z^{t}_{i}-z^{t}_{j})\cdot sgn(z^{s}_{i}-z^{s}_{j})=1\quad\forall i,j$ (14)
$\iff(z^{t}_{i}-z^{t}_{j})\cdot(z^{s}_{i}-z^{s}_{j})>0\quad\forall i,j$
$\iff z^{t}_{i}=F(z^{s}_{i})\quad\forall i\quad where\quad F^{\prime}(x)>0$
which means the optimal solution domain of ranking loss is achieved when there exists a monotonic increasing mapping between the student’s output and the teacher’s output. From the optimal solution domains of the aforementioned two, it is evident that the optimal solution domain for ranking loss is quite lenient and easily attainable, and it encompasses the optimal solution domain for KL divergence. During optimization with the differential form of rank loss, the closer it is to the optimal solution domain of ranking loss, the smaller the gradient provided for optimization becomes. This implies that during model optimization, the ranking loss will prevent the model from falling into the local optimum.
5 Experiment
Datasets and Baselines. In order to validate the efficacy and robustness of our proposed method, we conduct experiments on the CIFAR-100 [16] and ImageNet [24] datasets. CIFAR-100 is a significant dataset for image classification, comprising 100 categories, with 50,000 training images and 10,000 test images, each processed to a size of 32 $\times$ 32. ImageNet is a large-scale dataset utilized for image classification, comprising 1,000 categories, with approximately 1.28 million training images, and 50,000 test images, each processed to a size of 224 $\times$ 224. We apply our RKKD to a variety of logit-based Knowledge Distillation methods, including KD [12], CTKD [20], and MLKD [14]. We also compared various feature-based methods, including FitNet [1], CRD [30], and ReviewKD [4]. In addition, we also extend the proposed RKKD to the transformer student, which has a larger gap with teacher.
Implementation Details. Our experimental setup is primarily based on the settings of the compared baselines [12; 20; 42; 14]. For the CIFAR-100 dataset, we set the batch size to 64, and for the ImageNet dataset, we set the batch size to 512. We employ SGD [28] as the optimizer, with the number of epochs and learning rate settings consistent with the comparative baselines. The hyper-parameters $\alpha$, $\beta$ in Eq. 6 are set to be the same as the compared baselines to maintain fairness, and $\gamma$ are set equal to $\alpha$. We utilize 1 NVIDIA GeForce RTX 4090 to train models on CIFAR-100 and 4 NVIDIA GeForce RTX 4090 for training on ImageNet.
Table 1: CIFAR-100 Heterogeneous Architecture Results. The Top-1 Accuracy (%) is reported as the evaluation metric. The teacher and student have heterogeneous architectures. We incorporate the proposed ranking loss into the existing logit-based methods, with the performance gains indicated in parentheses..
Type Teacher ResNet32$\times$4 ResNet32$\times$4 ResNet32$\times$4 ResNet50 ResNet32$\times$4 WRN-40-2
79.42 79.42 79.42 79.34 79.42 75.61
Student SHN-V2 WRN-16-2 WRN-40-2 MN-V2 SHN-V1 SHN-V1
71.82 73.26 75.61 64.60 70.50 70.50
Feat FitNet [1] 73.54 74.70 77.69 63.16 73.59 73.73
CRD [30] 75.65 75.65 78.15 69.11 75.11 76.05
ReviewKD [4] 77.78 76.11 78.96 69.89 77.45 77.14
Logit KD [12] 74.45 74.9 77.7 67.35 74.07 74.83
KD+RKKD 76.58(+2.13) 75.18(+0.28) 78.50(+0.80) 70.45(+3.10) 75.98(+1.91) 76.13(+1.30)
CTKD [20] 75.37 74.57 77.66 68.67 74.48 75.61
CTKD+RKKD 77.01(+1.64) 75.71(+1.14) 78.61(+0.95) 70.18(+1.51) 76.67(+2.19) 76.80(+1.19)
DKD [42] 77.07 75.7 78.46 70.35 76.35 76.33
DKD+RKKD 77.60(+0.53) 75.99(+0.29) 78.75(+0.29) 70.90(+0.55) 77.36(+1.01) 76.43(+0.10)
MLKD [14] 78.44 76.52 79.26 71.04 77.18 77.44
MLKD+RKKD 78.53(+0.09) 76.83(+0.31) 79.86(+0.60) 71.66(+0.62) 77.63(+0.45) 77.87(+0.43)
Other Settings. In order to prevent the gradient explosions due to too small logit value used in RKKD, we only add normalization to handle the original logit in RKKD. Since using normalization does not affect the ranking, and we do not add normalization to the other loss terms, it does not affect our fairness as a plug-and-play function. More experiments are shown in the Appendix.
5.1 Main Result
CIFAR-100 Results. In our study, we conducted a comparative analysis of Knowledge Distillation (KD) outcomes across various teacher-student [9; 26; 39; 41; 13; 25] configurations as shown in Tab. 1 and Tab. 2. While Tab. 1 presents cases where the teacher and student models share the heterogeneous architecture, Tab. 2 illustrates instances of homogenous structures. Furthermore, as a plug-and-play method, we applied our RKKD in multiple logit-based distillation techniques. The incorporation of RKKD resulted in an average improvement of 1.83% in vanilla KD. In the context of existing state-of-the-art (SOTA) logits-based methods, both CTKD and MLKD exhibited significant gains. Notably, a subset of students surpassed their respective teachers after the distillation.
Table 2: CIFAR-100 Homogenous Architecture Results. The Top-1 Accuracy (%) is reported as the evaluation metric. The teacher and student have homogeneous architectures. We incorporate the proposed ranking loss into the existing logit-based methods, with the performance gains indicated in parentheses.
Type Teacher ResNet32$\times$4 VGG13 WRN-40-2 ResNet110
79.42 74.64 75.61 74.31
Student ResNet8$\times$4 VGG8 WRN-40-1 ResNet20
72.50 70.36 71.98 69.06
Feat FitNet [1] 73.50 71.02 72.24 68.99
CRD [30] 75.51 73.94 74.14 71.46
ReviewKD [4] 75.63 74.84 75.09 71.34
Logit KD [12] 73.33 72.98 73.54 70.67
KD+RKKD 74.74(+1.41) 74.14(+1.16) 74.49(+0.95) 71.09(+0.42)
CTKD [20] 73.39 73.52 73.93 70.99
CTKD+RKKD 75.59(+2.2) 74.76(+1.24) 74.86(+0.93) 71.08(+0.09)
DKD [42] 76.32 74.68 74.81 71.06
DKD+RKKD 76.61(+0.29) 75.1(+0.42) 74.94 (+0.13) 71.84(+0.78)
MLKD [14] 77.08 75.18 75.35 71.89
MLKD+RKKD 77.25(+0.17) 75.35(+0.17) 76.08(+0.73) 72.35(+0.46)
ImageNet Results. The results on ImageNet of KD in terms of top-1 and top-5 accuracy are compared in Tab.3. Our RKKD can achieve consistent improvement on the large-scale dataset as well.
Table 3: The top-1 and top-5 accuracy (%) on the ImageNet validation set. The teacher and student are ResNet50 and MN-V1
AT[38] OFD[10] CRD[30] KD[12] KD+RKKD
Top-1 69.56 71.25 71.37 70.50 71.54(+1.04)
Top-5 89.33 90.34 90.41 89.80 90.84(+1.04)
5.2 Extensions
KD for Transformer. Further exploration was conducted on the implications of rank loss when employing a transformer as both the teacher and the backbone. The experimental results indicate that the incorporation of RKKD yields significant improvements over the conventional knowledge distillation, as shown in Tab. 4. This denotes that RKKD is also applicable to distillation challenges predicated on transformer architectures. The implementation details of the Transformer experiments are attached in the Appendix.
Table 4: KD for Transformer. The Top-1 Accuracy (%) on the validation set of CIFAR-100. The Teacher model is ResNet56.
Student DeiT-Tiny [31] T2T-ViT-7 [37] PiT-Tiny [11] PVT-Tiny [33]
65.08 69.37 73.58 69.22
KD 71.11 71.72 74.03 72.46
KD+RKKD 73.25(+2.14) 72.49(+0.77) 74.33(+0.30) 73.42(+0.96)
KD for Object Detection. To further verify the generality of our proposed method in downstream tasks, we also conduct experiments under the setting of object detection. The results show that our proposed ranking loss can improve the performance of knowledge distillation method in object detection and achieve better performance than the same period of the feature-based object detection method, which is shown in Tab. 5.
Table 5: KD for Object Detection. All experiments are conducted on COCO2017 with the teacher as ResNet50 and the student as MobileNet-V2.
Method AP AP50 AP70
KD 30.13 50.28 31.35
FitNet[[1]] 30.20 49.80 31.69
FGFI[[32]] 31.16 50.68 32.92
KD+RKKD 31.99(+21.86) 53.80(+3.52) 33.37(+2.02)
Different Forms of Ranking Loss. In our investigation, we examined the performance of the three forms of distillation presented in Sec. 4.2, as depicted in Tab. 6. Form1 exhibited the most superior performance, followed by Form2, with Form3 trailing yet still enhancing the efficacy of the original knowledge distillation. This suggests that within the ranking loss, it is imperative to optimize considering the magnitude of the teacher’s logits differences as a coefficient for the loss, rather than merely optimizing as a sign function.
Table 6: Different Forms of Ranking Loss. The experiments are conducted on the CIFAR-100, with 9 heterogeneous and 7 homogeneous architectures. The average Top-1 accuracy (%) is reported.
Loss Form KD KD+Form1 KD+Form2 KD+Form3
Similar Structure 72.01 73.47(+1.46) 73.42(+1.41) 73.38(+1.37)
Different Structure 72.74 73.50(+0.76) 73.36(+0.62) 73.05(+0.31)
Ablation Study. We conducted extensive ablation studies to investigate the effectiveness of RKKD under various settings of $k$ and coefficients. Tab. 8 presents the distillation outcomes across different $k$ configurations. It is observed that a larger $k$ value yields superior performance, suggesting that the differential form of ranking loss more closely approximates the Kendall $\tau$ coefficient, thereby imparting stronger ranking knowledge. Tab. 7 delineates the performance of RKKD under varying coefficients. This setting effectively aligns the classification outcomes of the teacher and student models, thus significantly enhancing the distillation effect.
Table 7: Ablation of Weight. The Top-1 Accuracy (%) on the validation set of CIFAR-100.
Teacher ResNet32$\times$4 WRN-40-2 ResNet32$\times$4 ResNet50
79.42 75.61 79.42 79.34
Student ResNet8$\times$4 WRN-40-1 SHN-V2 MN-V2
72.50 71.98 71.28 64.60
KD(Baseline) 73.33 73.54 74.45 67.35
$\gamma=0.1$ 74.15 74.15 75.52 69.25
$\gamma=0.5$ 74.84 74.07 76.34 69.81
$\gamma=0.9$ 74.74 74.49 76.58 70.45
$\gamma=2$ 74.62 74.65 77.07 69.97
$\gamma=4$ 75.07 73.60 77.05 70.59
$\gamma=6$ 74.78 73.09 76.90 69.58
Table 8: Ablation of $k$. The Top-1 Accuracy (%) on the validation set of CIFAR-100.
Teacher ResNet32$\times$4 WRN-40-2 ResNet32$\times$4 ResNet50
79.42 75.61 79.42 79.34
Student ResNet8$\times$4 WRN-40-1 SHN-V2 MN-V2
72.50 71.98 71.28 64.60
KD(Baseline) 73.33 73.54 74.45 67.35
$k=0.1$ 73.13 73.75 75.47 68.9
$k=0.5$ 74.36 74.17 75.73 68.87
$k=1$ 74.74 74.49 76.58 70.45
$k=2$ 74.79 74.87 77.06 70.53
$k=4$ 75.56 74.48 77.21 70.81
$k=6$ 75.74 74.11 76.11 70.32
5.3 Analysis
Figure 3: Loss Landscape. The Left landscape shows the suboptimal solutions in the distillation task. After adding our ranking loss, the suboptimal solution is significantly reduced, as shown on the Right.
Visualization of Loss Landscape. To further investigate the role of ranking loss in the distillation process, we visualized the loss landscapes [19] of student models with and without the application of ranking loss during Knowledge Distillation (KD), as depicted in Fig. 4. It is evident that the student models distilled with ranking loss exhibit a markedly flatter loss landscape and fewer local optima compared to those without it. We hypothesize that during the optimization process, ranking loss can filter out certain local optima that, despite presenting a better overall loss performance, yield poorer classification outcomes. Consequently, ranking loss can effectively enhance the generalization performance of student models.
Figure 4: RKKD using Top-K and Min-K channels. Top-1 Accuracy (%) of Top-k/Min-k Ranking Loss. 0 and 100 on the x-axis represent KD and KD+Rank, respectively.
Top-k & Min-k Ranking. To further validate the beneficial knowledge present in smaller channels, we conducted comparative experiments using the top 10%, top 30%, top 50%, and min 10%, min 30%, min 50% channels. The experiments were performed across four combinations of homogeneous and heterogeneous teacher-student pairs, and the results are presented in Fig. 4. We observed that using the min-k channels achieves results similar to those obtained with top-k channels (as shown in Fig. 4(a),(b), and (c)). Additionally, in some cases, min-k channels provide even more beneficial information to aid student learning (as demonstrated in Fig. 4(d)), which indicates that the smaller channels also contain rich knowledge.
6 Conclusion
In this paper, we investigate the optimization process of logit distillation and identify that the Kullback-Leibler divergence tends to overlook the knowledge embedded in the smaller channels of the output. Moreover, KL divergence does not guarantee alignment between the classification results of the student and teacher models. To address this issue, we introduce a ranking loss that encourages the student model to pay more attention to the knowledge contained within the smaller channels, while also enforcing alignment with the teacher’s predictive outcomes. We provide a theoretical analysis demonstrating that ranking loss can serve as an effective auxiliary loss to KL loss. Building on the Kendall $\tau$ coefficient, we propose three differential loss forms tailored for the distillation context. Extensive experiments validate that our approach significantly enhances the distillation performance across various teacher-student architectures. As a plug-and-play auxiliary method, it also effectively improves the efficacy of logit-based distillation methods.
Appendix A Appendix / supplemental material
A.1 Limitations
We have not explored the effectiveness of our method on other downstream tasks.
A.2 KD for Transformer
We use the AdamW optimizer and train for 300 epochs with an initial learning rate of 5e-4 and a weight decay of 0.05. The minimum learning rate is 5e-6, and the patch size is 16. We set $\alpha$ = 1, $\beta$ = 1, $\gamma$ = 0.5, and batch size is 128. The GPU we used is a single RTX4090.
A.3 Ablation of Temperature
There has been extensive and detailed research on the temperature of the KL loss (e.g., CTKD[20], MLKD[14], LSKD[27]), achieving significant results. Unlike these studies, our approach aims to move away from focusing on the KL divergence and instead explore plug-and-play auxiliary losses to guide the KL divergence. Therefore, in our experiments, the temperature and other parameters of the KL divergence are kept the same as those in the respective baselines (with dynamic/multi-level temperatures used in CTKD+Ours and MLKD+Ours). The proposed ranking loss does not have a temperature parameter because the temperature does not affect the order of logits. Instead, the control over the distribution can be achieved by manipulating the steepness parameter $k$ of the sign function, and ablation experiments for $k$ are shown in Tab. 8. Meanwhile, although the temperature helps control the logit distribution, additional measures are needed to prevent the KL divergence from falling into suboptimal. We conducted an ablation study on the temperature, as shown in Tab. 9. The results demonstrate that ranking loss as a plug and play loss can bring further improvements under multiple temperature settings.
Table 9: Ablation of Temperature. All experiments are conducted on CIFAR-100 with the teacher as ResNet32×4 and the student as ResNet8×4.
Tempetature T = 4 T = 5 T = 6 T = 10
KD 73.33 73.39 73.43 73.55
KD+RKKD 73.56(+0.23) 73.49(+0.10) 74.36(+0.0.93) 74.04(+0.49)
A.4 Ablation of Normalization
Due to the random initialization of student, the output logits will be too variant at the start of optimization and occasionally lead to gradient explosions. Therefore, we add normalization to the ranking loss to stabilize the ranking loss optimization at the very beginning. We also supplemented a set of small-scale ablation experiments that showed that the improvement in ranking performance did not come from the normalization added to the ranking loss, as shown in Tab. 10 below:
Table 10: Ablation of Normalization+.
Teacher ResNet32$\times$4 WRN-40-2
79.42 75.61
Student SHN-V1 WRN-40-1
70.30 71.98
KD 74.07 73.54
KD+RKKD w/o Norm 76.38(+2.31) 74.07(+0.53)
KD+RKKD w Norm 75.98(+1.91) 74.49(+0.95)